A cylindrical aluminum block will be used in this experiment. If the block has a length 4cm, and a diameter 10.4cm, what is its volume in unit of cm3?
Answer #1

here,

radius of cylindrical, r = 10.4/2 = 5.2 cm

L = 4 cm

for the volume of the cylinder

volume of the cylinder = pi * r^2 * L

volume of the cylinder = pi * 5.2^2 * 4

volume of the cylinder = 339.7 cm^3

the volume of the cylinder is 339.7 cm^3
Answer #2

We can use the formula for the volume of a cylinder:

Volume=πr2h

Where:

  • r is the radius of the cylinder.

  • h is the height (or length) of the cylinder.

Calculation:

  1. Find the radius (r):
    The diameter is given as 10.4 cm, so the radius is half of that:

    r=Diameter2=10.4cm2=5.2cm

  2. Plug the values into the formula:

    Volume=π(5.2cm)2(4cm)

    • Length (h) = 4 cm

    • Radius (r) = 5.2 cm

  3. Calculate (5.2cm)2:

    5.2×5.2=27.04cm2

  4. Multiply by the length (h):

    27.04cm2×4cm=108.16cm3

  5. Multiply by π:
    Using π3.1416:

    Volume=3.1416×108.16cm3339.8cm3

Answer:

The volume of the cylindrical aluminum block is approximately 339.8 cm³.


answered by: Harshwardhan kunal